Shift report 20000831 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 31 2000 - 23:59:06 EDT

  • Next message: http account: "Shift report 20000901 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Aug 31 23:59:06 2000, Shift: 16:00-24:00
    Supervisor: YB, with: KH
    Log book#5 pages 66 - 69
    
    Shift Summary:
    ------------------------------------------------------------
    1. Injection studies most of the time. Beam is on since 9pm 
        Ib=10,000E9, Iy=17,900E9
    2. Run2488 173K trig1 vs. 152K collisions. Changed FFS angle
       to 8 degrees, runs 2489 and 2491 is still running.
    3. Couldn't turn on D2; called expert, then it happened 
       again. Now is OK. 
    4. BFS is stuck at 4 degrees. Several different support 
       came to look at the problem and finally had a short
       access (when RHIC personnel also needed an access).  
       It was determined that problem was beyond the scope
       of their expertise.  An expert will be called in and
       will work on it when it can be coordinated with RHIC
       for another access if that is needed.
    5. Implemented unpacking of T5 into BrRawDataInput.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 31 2000 - 23:59:45 EDT